# Search-based DNN Testing and Retraining with GAN-enhanced Simulations

### v1

@@ IV-E Threats to validity
To address internal validity threats, we carefully tested our implementation of DESIGNATE and its variants, and further relied on the original version of TACTIC provided by its authors.
@@
To address generalizability threats, we considered a state-of-the-art DNN implementing a computer vision task that is key for safety-critical systems (i.e., image segmentation). We therefore expect our results to be applicable across many industrial contexts, from automotive to robotics.
@@ VI Conclusion
Our empirical investigation, conducted using a state-of-the-art DNN for road segmentation, demonstrates the effectiveness of DESIGNATE in identifying failure-inducing inputs and maximizing input diversity. DESIGNATE consistently outperformed the baseline approaches, including the state-of-the-art TACTIC method, by generating inputs that led to worse DNN performance and more diverse failure scenarios. Furthermore, the realistic images generated by DESIGNATE proved to be valuable for DNN retraining, leading to improved performance on real-world images, thus outperforming alternative solutions and demonstrating that: (1) simulator images are not as effective as realistic images for DNN retraining (2) relying on GANs to generate images that simply change weather conditions is not as effective as combining GAN with a simulator to generate realistic images for DNN retraining. Based on our findings, we suggest that future work on DNN testing, instead of simply relying on simulators, should integrate GAN components transforming simulator outputs into more realistic ones.

### v2

@@ IV-E Threats to validity
To address internal validity threats, we carefully tested our implementation of DESIGNATE and its variants, and further relied on the original version of TACTIC and DeepJanus provided by their authors. ††margin: R3.7 Still related to internal validity, to prevent the generation of unrealistic images with GANs, our approach integrates a strategy to eliminate images that are likely to be unrealistic (i.e., images where the DNN performs better with the simulated image). Such heuristic seems effective since RQ2 results show that retraining the DNN with DESIGNATE images leads to improved results with a real-world test set. If DESIGNATE images were not realistic, the performance of the DNN would have shown limited or no improvement because it would have learned from images that are unrealistic, which is what happens with approaches without GAN. However, it is still possible that our filter may not be fully effective. In the future, we aim to extend our approach with additional methods such as histogram analysis [54] and other state-of-the-art approaches [59, 44] for the identification of out-of-domain and out-of-distribution images.
@@
To address generalizability threats, we considered two instances of a state-of-the-art DNN implementing a computer vision task that is key for safety-critical systems (i.e., image segmentation), for two very different environments (urban and Martian). We therefore expect our results to be applicable across many industrial contexts, from automotive to robotics.
@@ VI Conclusion
Our empirical investigation, conducted using a state-of-the-art DNN for road and martian terrains segmentation, demonstrates the effectiveness of DESIGNATE in identifying failure-inducing inputs and maximizing input diversity. DESIGNATE consistently outperformed the baseline approaches, including the state-of-the-art methods like TACTIC and DeepJanus, by generating inputs that led to worse DNN performance and more diverse failure scenarios. Furthermore, the realistic images generated by DESIGNATE proved to be valuable for DNN retraining, leading to improved performance on real-world images, thus outperforming alternative solutions and demonstrating that: (1) simulator images are not as effective as realistic images for DNN retraining (2) relying on GANs to generate images that simply change weather conditions is not as effective as combining GAN with a simulator to generate realistic images for DNN retraining. Based on our findings, we suggest that future work on DNN testing, instead of simply relying on simulators, should integrate GAN components transforming simulator outputs into more realistic ones.